piscrete Maths le fila 13. find f(2) and f(3) f. for (ins= 3/(n-1)+7
Answer #1

Answer:

f(2) = 3f(1) + 7 = 3*13 + 7 = 39 + 7 = 46

f(3) = 3f(2) + 7 = 3*46 + 7 = 138 + 7 = 145